# Computational Copyright: Towards A Royalty Model for AI Music Generation Platforms

### v1

@@ 2.1 Preliminary: The Concepts of Music Royalties
Prior to delving into the specific cases, it is essential to familiarize ourselves with the fundamental concepts and a few major types of music royalties that are prevalent in the industry.
@@
In order to understand the intricacies of the implementation of royalty models and their applicability to AI music generation platforms, we delve into case studies of two major platforms: Spotify and YouTube. Spotify is the largest music streaming platform in the world while YouTube is the largest video sharing platform. Both platforms have a significant amount of music content and generate revenue through multiple sources. Furthermore, despite various existing criticisms on these royalty models (Marshall, 2015; Trendacosta, 2020), they represent the status quo of how the current digital music market works. Therefore, designing royalty models for AI music generation platforms mimicking the ones for Spotify and YouTube would be a reasonable initial step in this area. In the following sections, we will examine the royalty models of these two platforms in detail.
2.2 Case Study: Spotify
Stakeholders.
Spotify’s royalty model involves several key groups of stakeholders, in addition to the streaming platform itself.
@@ 2.2 Case Study: Spotify
Stakeholders in groups 1, 2, and 3 receive revenue shares from Spotify, while groups 4 and 5 contribute to the generation of Spotify’s revenue. Typically, Spotify directly interacts with stakeholders in groups 2 and 3 — record labels, music publishers, music rights societies, and collecting agencies. Individual artists and creators often have contracts with these labels, publishers, or music rights agencies, and do not directly engage with Spotify in the financial aspect of their music streaming.
Revenue Sources.
The major revenue sources of Spotify can be divided into two categories: subscription and advertisement. In 2021, premium subscriptions accounted for 88% of Spotify’s revenue while advertisements accounted for the remaining 12% (Johnston, 2023). The two revenue sources lead to the formation of separate revenue pools, which are also calculated separately for different countries or regions.
Royalty Distribution.
Spotify employs a straightforward pro rata model to calculate the royalty distribution for each revenue pool. This model operates on the basis of streams, where a stream is defined as an individual music play exceeding a certain minimum duration. The calculation first counts the total number of streams on the whole platform for a specific period (usually a month), and then determines the share of streams for each artist or track as a percentage of the total streams. The royalty for each artist or track is calculated by applying their stream share to each revenue pool. This method ensures that royalty distribution is directly proportional to the popularity and streaming frequency of each artist’s or track’s work on the platform.
2.3 Case Study: YouTube
YouTube’s model for compensating music copyright owners is multifaceted, offering various methods for monetizing the content:
  1. 1.
    YouTube Partner Program: Music copyright owners can join the YouTube Partner Program, uploading music (videos) to their official channels. Revenue is shared based on user views of their content.
  2. 2.
    Content ID: Owners can earn from videos using their music through the Content ID system. This system uses fingerprinting and machine learning to identify copyrighted content in uploaded videos and allocates revenue from these videos to the copyright owners.
  3. 3.
    Licensing: Owners can also license their music directly to a YouTube video for a one-time payment.
The first method resembles Spotify’s royalty model. The second and the third methods are different as they involve a third-party video creator. While the first two methods can be considered forms of streaming royalties, the third method is more like a synchronization license.
Stakeholders.
Revenue Sources.
Royalty Distribution.
In the first two methods, YouTube first calculates the revenue share for each video in a manner similar to Spotify. However the calculation is more complex due to the diverse range of advertisement models and varying user interactions with the videos. For the second method, the revenue share for each video is further distributed to both video creators and music copyright owners, according to the Content ID system. This distribution can be automated, using an algorithm that assesses the actual use of music in the video, or it can be managed through manually negotiated agreements. In contrast, the third method employs a more straightforward royalty calculation, based directly on the licensing agreement.
Attributing Copyright Ownership.
A crucial aspect of YouTube’s royalty model is the challenge of attributing copyright ownership in the videos. Unlike Spotify, where content attribution is straightforward through stream counts, the incorporation of copyrighted music in the user-generated videos makes this task technically demanding. The Content ID system serves as the technical foundation that enables the second and third methods of YouTube’s revenue sharing. In the second method, it identifies music in videos and allocates revenue to the copyright owners. In the third method, while synchronization licenses might be obtained through third-party licensing platforms outside of YouTube, the presence of the Content ID system encourages them to secure these licenses. Although the system has its share of flaws and has faced criticism (Van der Sar, 2021; McKay, 2011; Trendacosta, 2020; Saadatpanah et al., 2020), including issues with false positives and negatives, it is still broadly embraced by the music industry.
2.4 Summary
A notable pattern of the royalty models on digital music/video platforms is that the payment is not directly made to each individual piece of music work [^3] The licensing model on YouTube, where video creators directly buy synchronization licenses for individual music, is an exception. However, note that this model is also enabled by the existence of the Content ID system.. Rather, the platforms accumulate overall revenue from their services involving a vast collection of music, and then distribute a portion of the revenue among the copyright owners.
The typical methods for determining how revenue is distributed can be summarized in two main steps:
  1. 1.
    Formation of Revenue Pools: Determine the total amount of revenue available for distribution in each revenue pool. These pools are defined by various criteria, including the sources of revenue and geographic regions.
  2. 2.
    Distribution Based on Access Frequency: Within each revenue pool, distribute the revenue proportional to the frequency that each piece of music is played or accessed.
For example, on Spotify, revenue pools are based on subscription and advertisement; the revenue distribution is determined by the stream counts. On YouTube, the step 2 needs to be done in two sub-steps, where one first determines the frequency of a video is played, then attributes the copyright ownership through the Content ID system.
3 Potential Royalty Models for AI Music Generation Platforms
3.1 The Business Models
While the landscape of AI music generation is still rapidly evolving, there have been a few common business models emerging (McFarland, 2023). We summarize these business models in terms of services, stakeholders, and revenue sources.
@@ 3.1 The Business Models
Stakeholders.
3.2 Potential Royalty Model Designs
Particularly, the business models for AI music generation platforms align with the pattern identified in Section 2.4, where the revenue is channeled through the platform, rather than directly compensating for each individual piece of music.
A potential royalty model for the AI music generation platforms can be structured in two main steps:
  1. 1.
    Formation of Revenue Pools: Similar to Spotify and YouTube, these platforms would first accumulate revenue, forming distinct pools based on different criteria such as revenue sources (subscriptions, advertisements, licensing fees).
  2. 2.
    Distribution Based on Access Frequency: The revenue from each pool would then be distributed based on the frequency at which each copyrighted work included in the training corpus is accessed during the service.
4 Attributing AI-Generated Music to Copyrighted Training Content
4.1 Formulation of the Attribution Problem
Attributing the influence of copyrighted training content on generated music essentially asks the question: “To what extent does each piece of music in the training corpus influence a specific piece of AI-generated music?”
To operationalize a royalty distribution mechanism, it’s crucial to establish a quantifiable definition of “influence”. This is a challenging open problem that can potentially be subjective. To this end, we suggest two perspectives to define the “influence”, one inspired by the machine learning literature and the other comes from the domain knowledge of music.
Perspective 1: Data Attribution.
The data attribution problem in machine learning refers to the task of identifying and quantifying the contribution of each data point in the training dataset to a given output of a machine learning model. Formally, this problem is often framed as follows: How does the removal of a particular data point from the training dataset and subsequent retraining of the model affect its output? This change in output serves as a measure of the removed data point’s influence on that specific model output (Koh and Liang, 2017). In the context of AI music generation, we can define the influence of a piece of training music on a piece of generated music in terms of the change in the likelihood of the model producing that generated music, assuming the model is retrained after removing the training music piece.
Perspective 2: Musical Influence.
4.2 Data Attribution for Music Generation
In this section, we first introduce formal definitions of AI music generation and the data attribution problem for music generation. Subsequently, we propose an algorithmic solution designed to quantitatively estimate the influence of each training piece on a particular piece of generated music.
4.2.1 AI Music Generation
In the field of AI music generation, there are two major paradigms: waveform music generation and symbolic music generation (Manzelli et al., 2018). Waveform music generation involves the direct synthesis of a music’s waveform, with examples including WaveNet (Oord et al., 2016). On the other hand, symbolic music generation involves creating music in a symbolic format, such as the Musical Instrument Digital Interface (MIDI) format. This paper focuses on symbolic music generation.
Symbolic Music Generation
We start by introducing the notations and the formal definitions of symbolic music generation. Symbolic music is characterized as a series of discrete events that collectively represent a segment of music. A music segment $m$ with $k$ events can be expressed as $(e_{1},e_{2},\ldots,e_{k})$. Here, each $e_{k}\in\mathcal{V}$ represents a specific event from a vocabulary set $\mathcal{S}$ that defines all the valid symbolic events. For instance, the vocabulary set might include events such as “Note 5 ON”, “Set velocity to 34”, or “Time shift 50ms”.
A symbolic music generation model, denoted as $h$, takes a prompt music segment and calculates the subsequent event’s conditional probability distribution as follows:
$P(e_{k}\mid e_{k-p},e_{k-p+1},\ldots,e_{k-1};h)=h(e_{k-p},e_{k-p+1},\ldots,e_{k-1}),$ (1)
where $P(e_{k}\mid e_{k-p},e_{k-p+1},\ldots,e_{k-1};h)$ is the conditional probability distribution of an event $e_{k}$, and $p$ indicates the length of prompt. This type of model formulation is known as an autoregressive model, which is one of the most popular model family for symbolic music. This approach sequentially predicts the probability of new events based on the preceding events, and then selects an event from these probabilities.
In this paper, we ground our study with Music Transformer (Huang et al., 2018), one of the most popular autoregressive symbolic music generation models that employs the Transformer architecture (Vaswani et al., 2017) for the generation model $h$. Suppose the size of the vocabulary set is $|\mathcal{V}|=V$. The model $h$ can be represented as a neural network with $V$ output units, each of which corresponds to a specific event in the vocabulary set, and the $V$ output units sum up to one. The output of the model corresponds to a probability distribution over the vocabulary set, and the model $h$ can be viewed as a classifier with $V$ classes. The model is trained by maximizing the log-likelihood over a training corpus of music.
4.2.2 Data Attribution for Symbolic Music Generation
Attribution Scores.
$I(m_{i},\hat{m})=f(\hat{m},h_{S\setminus\{m_{i}\}})-f(\hat{m},h_{S}),$ (2)
Two Levels of Attribution.
Estimating the Attribution Scores.
Directly calculating $I(m_{i},\hat{m})$ requires retraining a model for for each training data point $m_{i}$, which is computationally prohibitive. Fortunately, there has been a rich literature on efficient data attribution methods (Hammoudeh and Lowd, 2022), primarily designed for classification models. Furthermore, we shall see that these methods can be easily adapted to the autoregressive generative models like Music Transformer.
In particular, since generating one event can be viewed as a classification problem, we can directly apply existing data attribution methods for classification models to event-level attribution.
For the segment-level attribution, when the utility function is defined as the log-likelihood, i.e.,
$f(\hat{m},h)=\log P(\hat{m};h)$
we observe that, assuming $\hat{m}=(e_{k},\ldots,e_{l})$ for some $k<l$, then by Bayes’ rule,
$f(\hat{m},h)$ $=\log P(\hat{m};h)$
$=\sum_{i=k}^{l}\log P(e_{i}\mid e_{i-p},\ldots,e_{i-1};h)$
$=\sum_{i=k}^{l}f(\{e_{i}\},h).$
This implies that the attribution score $I(m_{i},\hat{m})$ can be decomposed as
$I(m_{i},\hat{m})$ $=f(\hat{m},h_{S\setminus\{m_{i}\}})-f(\hat{m},h_{S})$
$=\sum_{i=k}^{l}f(\{e_{i}\},h_{S\setminus\{m_{i}\}})-\sum_{j=k}^{l}f(\{e_{i}\},h_{S})$
$=\sum_{i=k}^{l}\left(f(\{e_{i}\},h_{S\setminus\{m_{i}\}})-f(\{e_{i}\},h_{S})\right)$
$=\sum_{i=k}^{l}I(m_{i},\{e_{i}\}),$
There are several off-the-shelf data attribution methods that can be applied to estimate the attribution scores for the autoregressive symbolic music generation models, such as Influence function (Koh and Liang, 2017), TraceIn (Pruthi et al., 2020), and TRAK (Park et al., 2023). In our study, we choose to use TRAK for its efficiency and state-of-the-art performance. We denote the estimated attribution score for $m_{i}$ on $\hat{m}$ as $\hat{I}(m_{i},\hat{m})$.
4.3 Experimental Evaluation
In this section, we empirically evaluate the proposed event-level attribution and segment-level attribution methods.
4.3.1 Experimental Setup
@@
We use the MIDI and Audio Edited for Synchronous TRacks and Organization (MAESTRO) dataset (v2.0.0) (Hawthorne et al., 2019) for our experiments. The dataset contains more than 200 hours of piano performances and is stored in MIDI format. The MIDI format is a widely recognized industry standard for representing musical information. Generally, it encodes music as a sequence of MIDI messages. Each type of MIDI message carries different additional attributes. For example, “Note ON” MIDI messages include attributes like velocity, duration, and pitch key. We refer readers to a tutorial (de Oliveira and de Oliveira, 2017) for more details about the MIDI format.
Following the setting of Music Transformer (Huang et al., 2018), we define a vocabulary set of size 388, which includes “NOTE ON” and “NOTE OFF” events for 128 different pitches (256 events in total), 100 “TIME SHIFT” events for different duration lengths, and 32 “VELOCITY” events that set the velocity of the next “NOTE ON” event. The raw data is preprocessed as sequences of events from this vocabulary set.
@@
For the implementation of attribution score calculation, we utilize the TRAK Park et al. (2023) toolkit for both event-level and segment-level attribution. We use the default values in TRAK toolkit for all the hyperparameters. For the ensemble setting, we use 50 models that are independently trained on 50% random subset of the training set.
4.3.2 Evaluation Metrics
We introduce evaluation metrics formalizing the two perspectives of “influence” discussed in Section 4.1.
Retraining Rank Correlation.
To evaluate the estimated attribution scores from the data attribution perspective, we can compare them with the ground truth attribution scores defined in Eq. (2). In the data attribution literature (Koh and Liang, 2017; Ilyas et al., 2022), the comparison is typically measured by Spearman’s rank correlation. Formally, for a training dataset with $n$ data points, one will calculate the rank correlation between $\{\hat{I}(m_{i},\hat{m})\}_{i=1}^{n}$ and $\{I(m_{i},\hat{m})\}_{i=1}^{n}$.
Musical Similarity.
For the musical influence perspective, there are multiple aspects mentioned in Section 4.1. In our study, we focus on the similarity of musical styles. A common approach to quantitatively evaluate musical style similarity is by extracting features from the music (Slaney et al., 2008). In this study, we identify three features used in Spotify APIs  [^4] https://developer.spotify.com/documentation/web-api/ to characterize a piece of music.
  1. 1.
    Loudness: It measures the overall velocity of a music segment. We define it as the average velocity of the “Note ON” events within the segment.
  2. 2.
    Key: It measures the average pitch height of all “Note ON” events in the music segment.
  3. 3.
    Duration: It measures the total length of the music segment in time, calculated as the sum of the time deltas of all events.
4.3.3 Experimental Results
Retraining Rank Correlation.
We form the set $U$ with 100 random subsets, each contains 50% of the training samples. We calculate the rank correlations on 178 generated music and report the average rank correlations for different data attribution methods in Table 1. “Random score” refers to assigning a random score for each training sample. “TRAK (segment-level)” and “TRAK (event-level)” refers to different levels of attribution. In comprison to the random baseline, both levels of attribution methods have achieved significantly positive correlations with the ground-truth scores. This indicates that there exist computationally feasible solutions that can reasonably attribute the generated music to the copyrighted training content, thus solving the key technical bottleneck for establishing a royalty model. In addition, we observe that event-level attribution seems to be easier than segment-level attribution. This leads to an interesting question about the proper granularity of attributing generated music, which we leave for future exploration.
Table 1: The average Spearman’s rank correlation among 178 generated music for different data attribution methods.
Data attribution method Random score TRAK (segment-level) TRAK (event-level)
Spearman’s rank correlation 0.0091 0.301 0.359
Musical Similarity.
For each generated music, we order the training music pieces by their attribution scores. Figure 1 shows the results of musical similarity between the generated music and training music pieces being decreasingly influential on the generated music. We observe a clearly decreasing trend of similarity in these plots, which suggests that the data attribution methods also capture some influence in terms of musical styles.
Figure 1: We calculated 178 pairs of generated music and the training sample with different ranks of the attribution score’s loudness, key, and duration. The boxplots represent the statistics of the Pearson correlation coefficient for different rank ranges and are plotted for each feature.
5 Conclusion
Acknowledgement
The authors would like to thank Yi Liang for helpful discussions on evaluating musical styles.

### v3

@@ Appendix A A Primer on the Concepts of Music Royalties
It is essential to familiarize ourselves with the fundamental concepts and a few major types of music royalties that are prevalent in the industry.
@@
Appendix B Detailed Stakeholder Description
@@
Appendix C AI Music Generation Platforms Business Models
@@
Appendix D More Details of Experimental Setup
@@
We use the MIDI and Audio Edited for Synchronous TRacks and Organization (MAESTRO) dataset (v2.0.0) Hawthorne et al. [2019] for our experiments. The dataset contains more than 200 hours of piano performances and is stored in MIDI format. The MIDI format is a widely recognized industry standard for representing musical information. Generally, it encodes music as a sequence of MIDI messages. MIDI messages include attributes like velocity, duration, and pitch key. We refer readers to a tutorial de Oliveira and de Oliveira [2017] for more details about the MIDI format.
Following the setting of Music Transformer Huang et al. [2018], we define a vocabulary set of size 388, which includes “NOTE ON” and “NOTE OFF” events for 128 different pitches (256 events in total), 100 “TIME SHIFT” events for different duration lengths, and 32 “VELOCITY” events that set the velocity of the following “NOTE ON” events. The raw data is preprocessed as sequences of events from this vocabulary set.
@@
For the implementation of attribution score calculation with TRAK Park et al. [2023], we utilize the TRAK toolkit for both event-level and segment-level attribution. We use the default values in TRAK toolkit for all the hyperparameters. For the ensemble setting, we use 50 models that are independently trained on 50% random subset of the training set. For TracIN Pruthi et al. [2020], we utilize the approximation version on the last linear layer, i.e., $(\nabla_{y}f(y)\cdot\nabla_{y^{\prime}}f^{\prime}(y^{\prime}))(x\cdot x^{\prime})$. $x$ and $x^{\prime}$ represent the last second layer representation of training and test samples. $y$ and $y^{\prime}$ represent the model output of training and test samples.
Appendix E Discussion on Musical Similarity and Attribution Scores
Our experiments reveal a synergy between musical similarity and attribution scores, supporting the intuitive belief that training music pieces more akin to generated ones should receive greater attribution. This section discusses the distinctions between musical similarity and attribution scores, illustrating why relying solely on similarity assessments might be inadequate for addressing the challenge of attribution.
Initially, the concept of musical similarity is model-independent, indicating that it consistently assigns similar attributes regardless of the model used. This characteristic is at odds with the fact that different models, even when trained on the same dataset, can exhibit varied behaviors. Furthermore, musical similarity fails to account for the “interactions” within the training dataset. For instance, the presence of multiple similar music pieces in the dataset can influence the contribution of each piece.
Appendix F Experiment Settings and Further discussion for Robustness
Duplication.
For the experiment of the adversarial attack using duplication, we add multiple duplicate copies of a training sample into the dataset and then recalculate the attribution scores. The experimental setup, including dataset preparation, model training, generation processes, and data attribution methods, remains consistent with the procedures outlined in Appendix D. We select three independent training music sequences with length of 256 events as the target samples for duplication to assess the sensitivity to the target samples. We test varying quantities of duplicated copies—1, 2, 5, 10, and 100—to explore how changes in the number of copies affect the attribution scores.
This phenomenon that any number of duplicated copies would bring the scores to nearly zero can be understood through the lens of data attribution definition, particularly the concept of leave-one-out influence as described in Eq. 1. The minimal influence observed can be attributed to the presence of at least one duplicated sample in the leave-one-out dataset, thereby diminishing its overall impact.
Copy-and-Replace.
For the experiment of the adversarial attack using copy-and-replace, we copy a segment from a training sample (source) that has top attribution scores and replace another training sample (target) at the same position with it. The experimental setup, including dataset preparation, model training, generation processes, and data attribution methods, remains consistent with the procedures outlined in Appendix D. We select two independent training music sequences with top attribution scores as the source samples to be copied to assess the sensitivity to the source samples. We test varying quantities of copied events—1, 2, 4, 8, 16, and 32—to explore how changes in the number of copied events affect the attribution scores.